OpenStudy (anonymous):

What is the equation of the circle with the center at the origin and with a radius of 9?

OpenStudy (anonymous):

The equation of a circle is \[(x-a)^2 + (y=b)^2 = r^2\] where (a,b) is the center and r is the radius. So substitute your values in :)

OpenStudy (aravindg):

Hi there welcome to openstudy !